Operation Summary for October 17, 2011 - October 19, 2011 |
Monday October 17 The day shift began with Operations monitoring protons in the Recycler (1.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, and with the NuMI and MiniBooNE experiments taking beam. During the day shift, Linac personnel began a 10 MeV study at 8:55 AM; this study interrupted beam at times. Operations established beam to the Meson FTBF experiment T-992 at 10:25 AM. At 11:15 AM, operators reported that a Booster RF station (BRF14) tripped off for the first of four times in about three hours due to sparking. At 2:49 PM, Operations turned off beam to MI for three minutes so that a Power Supply person could work on a Lambertson (LAM42). Operators loaded protons into the Antiproton Source (Pbar) Accumulator at 3:40 PM for a study. During the evening shift, A Mechanical Support engineer reported at 4:35 PM that he had raised the MI-62 LCW supply temperature from 85˚ F to 90˚ F; this was done to try and resolve the NuMI magnet (HV101) trips on high temperature. The Pbar Accumulator study ended at 6:17 PM. At 8:42 PM, BRF7 tripped off; it reset. Operations disabled beam to T-992 at 10 PM. Operators conducted a Booster injection tune at 11:03 PM. At 11:20 PM, BRF12 tripped off; it reset. Tuesday October 18 The midnight shift began with Operations monitoring protons in the Recycler (2.9E10), and with the NuMI and MiniBooNE experiments taking beam. During the midnight shift, at 1:57 AM, MIRF18 tripped off for the first of seven times in three and a half hours; it reset. Operators adjusted the station’s feed forward setting. At 3:56 AM, BRF7 tripped off for the first of twelve times in about an hour and a half; it reset. Operators adjusted the station’s modulator voltage. At 6:56 AM, BRF7 tripped off three times within a minute; it reset. Neutron Therapy Facility (NTF) personnel began at 7:35 AM preparing for treating a patient later in the morning. A RF person began work on BRF7 at 7:57 AM. During the day shift, at 8:26 AM, LAM42 tripped off; it reset. Operations established beam to T-992 at 10:01 AM. Pbar personnel began a study at noon. LAM42 tripped off at 1:49 PM; it reset. MI personnel began a bunch rotation study at 2:34 PM; this study lowered NuMI’s beam intensity. Pbar personnel ended their study at 3:47 PM. Linac and Booster personnel also conducted studies during this shift. During the evening shift, Booster personnel completed their study at 5:12 PM. At 5:45 PM, LAM42 tripped off for the first of four times in about four hours; it reset. At 8:26 PM, a Linac Klystron RF cavity (KRF5) suffered a vacuum burst; BRF14 and 15 tripped off due to bad beam positions caused by the burst; everything reset. Operations disabled beam to T-992 at 8:30 PM; this ended their current run at Fermilab. Wednesday October 19 The midnight shift began with Operations sending beam to the NuMI and MiniBooNE experiments. During the midnight shift, at 1:13 AM, operators prepared Switchyard for a scheduled access later in the morning. Operations turned off the MiniBooNE beam at 6:05 AM for an access that will occur later in the morning. At 6:17 AM, Operators prepared the Meson enclosure M03 for access later in the morning. Operations Coordinator (OPSCO) Plans for Today The plans for today are to end all machine access by noon and then resume normally operations. |
|
More Information Comments and Suggestions |
| last modified 10/19/2011 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |